# Locus: Agentic Predicate Synthesis for Directed Fuzzing

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
Directed Greybox Fuzzing (DGF) aims to search for program inputs leading its execution to reach specific target program states, e.g., an index variable used to access an array is out of the array bounds, thereby uncovering potential bugs or vulnerabilities. It is widely used in software engineering and security applications, including debugging system crashes(Tan et al., 2023; Wang et al., 2021), testing patches(Böhme et al., 2013; Marinescu and Cadar, 2013), verifying bug reports from static analysis(Christakis et al., 2016; Murali et al., 2024), and generating Proof-of-Concept (PoC) exploits of vulnerabilities(Brumley et al., 2008; Liu et al., 2024a). While the original purpose of directed fuzzing is largely not for discovering new bugs, i.e., it needs a specified target, it has found numerous impactful zero-day vulnerabilities (Liang et al., 2020; Nguyen et al., 2020; Wang et al., 2025; Cao et al., 2023; Shah et al., 2022).
@@
DGF is challenging, as the target program states are often deeply nested in the program, while the search space introduced by the complexity of real-world software is prohibitively large. To speed up the search and schedule promising inputs, most existing techniques leverage the control flow proximity, e.g., the distance to target location in the control flow graph, or heuristics based on the semantics of the branch predicates (Chen and Chen, 2018; Li et al., 2017; Shah et al., 2022), e.g., the target state if x==42 has the distance metric of |x42||x-42|. However, such feedback is sometimes too sparse or indirect to reliably measure the progress, especially when there is a long chain of implicit preconditions guarding the target program states (Mitropoulos et al., 2023; Holler et al., 2012; Dutra et al., 2023; Rigger and Su, 2020; Klees et al., 2018; Zhang et al., 2023; Kim et al., 2023; Sherman and Nagy, 2025; Aschermann et al., 2020). For example, triggering CVE-2018-13785 in libpng requires a PNG file to satisfy a precise sequence of preconditions, i.e., valid signature, correct chunk layout, specific IHDR fields (e.g., bit depth, color type), and a magic image width (0x55555555), to trigger an integer overflow (Kim et al., 2023), while the predicates to explicitly check these conditions are largely absent in the code to provide an incremental progress guidance.
@@
To capture the intricate feedback to improve search efficiency, more advanced approaches identify progress-capturing constraints in the program to drive execution towards satisfying specific temporal orders and preconditions (Meng et al., 2022; Holler et al., 2012; Lee et al., 2021; Liang et al., 2020; Li et al., 2017; Fioraldi et al., 2021; Padhye et al., 2019; Ispoglou et al., 2020; Balzarotti, 2021; Aschermann et al., 2020). However, these constraints are often manually crafted by experts and tailored to a few specific target state types, e.g., focusing on a temporal memory safety bug like a use-after-free by enforcing an allocate–free–use sequence. As different programs can have diverse target states and disparate functionalities, the feedback metric effective in one case may not generalize to another (Chen et al., 2019).
@@
As Machine Learning (ML) and Large Language Models (LLMs) have demonstrated surprising code reasoning capabilities, there has been a growing interest in extending such capabilities to help guide fuzzing (Patra and Pradel, 2016). A common approach is to employ LLMs to directly generate inputs or grammar-aware input generators (i.e., fuzz driver or harness) (Liu et al., 2024a; Huang et al., 2024b; Liu et al., 2024c; Shi et al., 2024; Zhang et al., 2024d; Lyu et al., 2024; Chen et al., 2025; Yang et al., 2025a), where the preconditions useful to reach target states are expressed as part of the input grammar constraints. However, not all conditions to reach target states can be easily represented as input grammars. For example, Figure 1 shows that an intermediate state (!found_plte) necessary to trigger the buffer overflow in libpng only emerges in the middle of the execution and cannot be easily checked at the input level. More importantly, such a task formulation is particularly challenging for LLMs, as reasoning from the target program states all the way to the input often requires an extremely long context and convoluted reasoning chain (Jiang et al., 2024; Li et al., 2025; Ding et al., 2024a; Li et al., 2024a; Risse et al., 2025; Risse and Böhme, 2024; Li et al., 2024b; Risse et al., 2025), a common pitfall for LLM hallucination (Sahoo et al., 2024; Yao et al., 2025), let alone the challenge to verify the correctness of the LLM generated inputs and harnesses, which can significantly impede the fuzzing progress if the generated constraints are incorrect (Padhye et al., 2019; Ispoglou et al., 2020).
@@
Our approach  We present Locus, a new framework that integrates LLMs’ code reasoning capabilities for directed fuzzing by synthesizing semantically meaningful and verifiable predicates to guide the search. Unlike existing LLM-based approaches that focus on constraining the search space at the input (harness) level, which poses a high reasoning burden on LLMs and can be hard to verify, Locus generalizes the constraint generation to be at arbitrary program points. Specifically, given a target state to reach, Locus automates the reasoning about the intermediate program states that verifiably relax the target states and synthesizes predicates as a curriculum to capture the gradual progress towards reaching them. Such predicates serve as the preconditions dominating all executions to reach the target states, providing fine-grained progress feedback to guide the fuzzers for input scheduling and early termination(Luo et al., 2023; Huang et al., 2022; Srivastava et al., 2022a). As these predicates are implemented as source-level instrumentation, they are agnostic to any fuzzer implementations and can thus be integrated without any customization. Moreover, as the instrumentation is a one-time offline process, the cost of running Locus can be amortized in all the succeeding fuzzing procedures.
@@
Agentic design  Automating the generation of effective progress-capturing predicates for diverse target states involves nontrivial reasoning across the entire codebase, spanning multiple functions and their associated data and control flows. Simple prompting techniques, such as in-context learning, chain-of-thought (CoT) prompting, retrieval-augmented generation, etc., can hardly support sophisticated reasoning at the repository level. To this end, Locus features an agentic synthesizer-validator workflow, equipped with diverse program analysis tools to support the traversal of the control flow graph, tracking data dependencies, retrieving function calls, and symbolic execution. They serve as the agent’s action space, allowing the LLM to reason via CoT and act by calling them (Yao et al., 2023) to iteratively propose, refine, and validate candidate predicates.
@@ 2.1. Directed Fuzzing
Canary  Directed fuzzing aims to generate inputs that drive the program execution to reach predefined program states. In this paper, we rely on the concept canaries (Hazimeh et al., 2020) to explicitly represent the target states, which are considered reached when the corresponding canary condition is satisfied. Early works like AFLGo(Böhme et al., 2017) and Beacon(Huang et al., 2022) treat canaries primarily as the reachability to particular program points, i.e., specific line numbers in a code file. Recent works (Hazimeh et al., 2020; Weissberg et al., 2024) adopt assertion-based canaries, where predicates are explicitly inserted to check vulnerability-triggering conditions at runtime. An example of a canary is highlighted in the bottom-left code snippet of Figure 1.
@@ 2.2. Motivating Example
A simplified CFG of this vulnerability is shown in Figure 1(a), where all the nodes highlighted in gray are equidistant to the canary node (highlighted in yellow), but only one node PLTE is in the path that can lead to the vulnerability (annotated as green arrows). Based on this, fuzzers can only perform conservative pruning (annotated as dashed arrows) while omitting a large part of the paths that are irrelevant to the vulnerability, e.g., all solid black arrows. Capturing the progress towards reaching this canary requires specialized knowledge about libpng parsing logic. However, developing such specialized progress-capturing state representations requires manual efforts from experts and cannot easily generalize to different vulnerabilities and programs.
@@
Figure 1(b) shows an example of the input generator synthesized by LLM-based approaches that enforced some specific input constraints. However, this approach is inherently limited as program inputs can have sophisticated structures that cannot be easily enforced at the input level, e.g., the PNG file contains a compressed data chunk IDAT. Therefore, the constraints posed on the input generated by the LLM often reduce to generic input grammars, e.g., png_sig_cmp only trivially ensures the input is a valid PNG file. We also note that the synthesized generators cannot be easily checked to determine whether they can effectively help reach the canary. For example, png_write_frame_head constrains the generated input to a special PNG type APNG that is essentially impossible to reach the canary, but automatically checking this fact is challenging. Moreover, to constrain the input to satisfy certain properties necessary to reach the target, the generator sometimes needs to repeat the input processing logic presented in the execution path. Such repeated execution introduces additional overhead.
@@
Existing fuzzers struggle with such a vulnerability due to the complex parsing logic. Since PNG chunks may appear in arbitrary order and many are optional, the primary parsing routine implements a loop in png_read_info to iteratively process each chunk. Within this loop, multiple branches exist, and each is responsible for handling a specific chunk type, e.g., IHDR, IDAT, PLTE. This loop makes it possible to identify whether a given PNG file contains a PLTE chunk. However, since these branches are syntactically parallel and executed without a fixed order, they all appear equidistant from the vulnerability site (png_set_PLTE) in the control flow graph. As a result, the naive path distances cannot distinguish between them to prioritize one path over the other. This explains why, in Table 2 and Table 3, all existing fuzzers take a substantial amount of time to reach this target state in libpng.
Locus starts with generating a predicate to check whether the input PNG file contains a PLTE chunk at the caller of the canary function ( 1). This implies that if the PNG file does not contain a PLTE chunk, the execution can terminate, as it is impossible to reach the target state. While this predicate is semantically correct, i.e., the symbolic execution confirms there is no feasible path to satisfy the target state while violate the generated predicate (Section 3.4), and can help filter out non-palette-based PNG files, it is almost redundant as there is an existing one immediately after the generated check (if (color_type & MASK_PALETTE)), and thus the generated predicate could barely help the fuzzer.
To address this issue, Locus includes another refinement iteration (Algorithm 1) to propagate this predicate to a location closer to the program entry such that the infeasible execution can be terminated earlier. By traversing the call graph and retrieving the functions along the call chain, the refinement proposes a new location in function png_read_info right after the parsing loop ( 2). Such a propagated predicates will also be validated again to ensure it remains a strict precondition to reach the target states. At the next iteration of refinement, Locus proposes to propagate this predicate further inside the branch that parses the IDAT chunk ( 3). That is because the specification of the PNG file requires that the optional PLTE chunk must appear before the IDAT chunk. Therefore, by the time the parsing procedure reaches the IDAT chunk, we can already determine whether the input PNG file contains a PLTE chunk.
@@ 3.1. Formalization
We next illustrate how our agentic synthesizer-validator workflow can produce an admissible instrumented program $P^{\prime}$ which helps provide rich semantic feedback for the directed fuzzer to reach vulnerable states $S_{v}$.
@@ 3.2. Agent Toolset
Existing agentic practices for software analyses, such as bug fixing or fault localization, are often equipped with lightweight command-line tools to perform local reasoning(Paul Gauthier, 2024; Zhang et al., 2024b; Yang et al., 2024b). This is effective because the root cause of a bug is often spatially close to the observable failure, allowing the agent to retrieve relevant context from a narrow portion of the entire codebase. However, in our task formulation, predicates can be synthesized at any program point. As a result, local reasoning is insufficient, and the agent needs additional tools to support analysis of a broader code context.
To equip the agent with long-context code reasoning at arbitrary program points, we provide a suite of tools that Locus can invoke to navigate the entire codebase and retrieve relevant code snippets. Particularly, in addition to common tools like code search and file listing, Locus integrates specialized tools for traversing program graphs, including call graphs and reference graphs. Through the call graph APIs, the synthesizer agent can retrieve function call relationships and reason the interprocedural control flow, while the reference graph API allows it to identify variable usages, pointer dereferences, and data access patterns across the codebase. Table 1 shows the complete toolset and the ratio of their actual usage in our experiments. Among them, graph traversal accounts for over a quarter of all API invocations.
It is important to note that these graphs are only provided as supplementary references to Locus, as they are all derived statically and may miss program behaviors such as dynamic dispatches and indirect calls. We observe that the LLM used in Locus is capable of leveraging its learned knowledge to bridge this gap. For example, in the case of vulnerability TIF002 (see Table 2 and Table 3), Locus successfully resolved the indirect function pointer tif->tif_decoderow to its concrete implementation PixarLogDecode.
@@
2:an target-conditional equivalent program $P^{\prime}$
@@ 3.3. Synthesis
Algorithm 1 elaborates on the workflow of Locus. For each candidate predicate, it iteratively localizes and generates the predicates (lines 4–13), which are then refined and validated for both syntax and semantics (line 8). Once validated, the predicates will instrument the original program to produce a fuzzing-admissible program (line 16).
@@
For each approximated characteristic, the synthesizer needs to identify an appropriate program point $l$ where the predicate can be expressed in terms of the variables and expressions in scope. Constraining LLM to directly identify the exact program point is often too challenging and unreliable, so the initial stage only asks the synthesizer to select a candidate function rather than a precise program point. Given the reasoning context generated by the LLM so far, the synthesizer attempts to generate an initial predicate $\phi_{l}$ in the candidate function. Once $\phi_{l}$ passes the validation stage (Section 3.4), we refine it in the following iterations. The goal of such a refinement is to identify an earlier program point where the same semantic check can be performed. This allows the program to reject invalid inputs sooner in the execution, enabling the fuzzer to explore more valid mutations within the same time budget.

### v3

@@ Abstract.
Directed Fuzzing, Test Generation, LLM Agent
journalyear: 2026copyright: ccconference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 2026 IEEE/ACM 48th International Conference on Software Engineering (ICSE ’26), April 12–18, 2026, Rio de Janeiro, Brazildoi: 10.1145/3744916.3773102isbn: 979-8-4007-2025-3/2026/04ccs: Security and privacy Software and application securityccs: Computing methodologies Machine learning
@@ 1. Introduction
Directed Grey-box Fuzzing (DGF) aims to search for program inputs leading its execution to reach specific target program states, e.g., an index variable used to access an array is out of the array bounds, thereby uncovering potential bugs or vulnerabilities. It is widely used in software engineering and security applications, including debugging system crashes(Tan et al., 2023; Wang et al., 2021), testing patches(Böhme et al., 2013; Marinescu and Cadar, 2013), verifying bug reports from static analysis(Christakis et al., 2016; Murali et al., 2024), and generating Proof-of-Concept (PoC) exploits of vulnerabilities(Brumley et al., 2008; Liu et al., 2024a). While the original purpose of directed fuzzing is largely not for discovering new bugs, i.e., it needs a specified target, it has found numerous impactful zero-day vulnerabilities (Liang et al., 2020; Nguyen et al., 2020; Wang et al., 2025; Cao et al., 2023; Shah et al., 2022).
@@
DGF is challenging, as the target program states are often deeply nested in the program, while the search space introduced by the complexity of real-world software is prohibitively large. To speed up the search and schedule promising inputs, most existing techniques rely on metrics based on control flow proximity, e.g., the distance to target location in the control flow graph, or heuristics based on the semantics of the branch predicates(Chen and Chen, 2018; Li et al., 2017; Shah et al., 2022), e.g., the target state if x==42 has the distance metric of |x42||x-42|. However, such feedback is sometimes too sparse or indirect to reliably measure the progress, especially when there is a long chain of implicit preconditions guarding the target program states (Mitropoulos et al., 2023; Holler et al., 2012; Dutra et al., 2023; Rigger and Su, 2020; Klees et al., 2018; Zhang et al., 2023; Kim et al., 2023; Sherman and Nagy, 2025; Aschermann et al., 2020). For example, triggering CVE-2018-13785 in libpng requires a PNG file to satisfy a precise sequence of preconditions, i.e., valid signature, correct chunk layout, specific IHDR fields (e.g., bit depth, color type), and a magic image width (0x55555555), to trigger an integer overflow (Kim et al., 2023), while the predicates to explicitly check these conditions are largely absent in the code to provide an incremental progress guidance.
@@
To capture the intricate feedback to improve search efficiency, more advanced approaches identify progress-capturing constraints in the program to drive execution towards satisfying specific temporal orders and preconditions (Meng et al., 2022; Holler et al., 2012; Lee et al., 2021; Liang et al., 2020; Li et al., 2017; Fioraldi et al., 2021; Padhye et al., 2019; Ispoglou et al., 2020; Balzarotti, 2021; Aschermann et al., 2020). However, these constraints are often manually crafted by experts and tailored to specific target state types, e.g., focusing on a temporal memory safety bug like a use-after-free by enforcing an allocate–free–use sequence. As different programs can have diverse target states and disparate functionalities, the feedback metric effective in one case may not generalize to another (Chen et al., 2019).
@@
As Machine Learning (ML) and Large Language Models (LLMs) have demonstrated surprising code reasoning capabilities, there has been a growing interest in extending such capabilities to help guide fuzzing (Patra and Pradel, 2016). A common approach is to employ LLMs to directly generate inputs or grammar-aware input generators (i.e., fuzz driver or harness) (Liu et al., 2024a; Huang et al., 2024b; Liu et al., 2024c; Shi et al., 2024; Zhang et al., 2024d; Lyu et al., 2024; Chen et al., 2025; Yang et al., 2025a), where the preconditions for reaching target states are expressed as part of the input grammar constraints. However, not all conditions to reach target states can be easily represented as input grammars. For example, Figure 1 shows that an intermediate state (!found_plte) necessary to trigger the buffer overflow in libpng only emerges in the middle of the execution and cannot be easily checked at the input level. More importantly, such a task formulation is particularly challenging for LLMs, as reasoning from the target program states all the way to the input often requires an extremely long context and convoluted analysis chain (Jiang et al., 2024; Li et al., 2025; Ding et al., 2024a; Li et al., 2024a; Risse et al., 2025; Risse and Böhme, 2024; Li et al., 2024b; Risse et al., 2025), a common pitfall for LLM hallucination (Sahoo et al., 2024; Yao et al., 2025), let alone the challenge to verify the correctness of the LLM generated inputs and harnesses, which can significantly impede the fuzzing progress if the generated constraints are incorrect (Padhye et al., 2019; Ispoglou et al., 2020).
@@
Our approach  We present Locus, a new framework that integrates LLMs’ code reasoning capabilities for directed fuzzing by synthesizing semantically meaningful and verifiable predicates to guide the search. Unlike existing LLM-based approaches that focus on constraining the search space at the input (harness) level, which poses a high reasoning burden on LLMs and can be hard to verify, Locus generalizes the constraint generation to be at arbitrary program points. Specifically, given a target state to reach, Locus automates the analysis about the intermediate program states that verifiably relax the target states and synthesizes predicates as a curriculum to capture the gradual progress towards reaching them. Such predicates serve as the preconditions dominating all executions to reach the target states, providing fine-grained progress feedback to guide the fuzzers for input scheduling and early termination(Luo et al., 2023; Huang et al., 2022; Srivastava et al., 2022a). As these predicates are implemented as source-level instrumentation, they are agnostic to any fuzzer implementations and can thus be integrated without any customization. Moreover, as the instrumentation is a one-time offline process, the cost of running Locus can be amortized in all the succeeding fuzzing campaigns.
@@
Agentic design  Automating the generation of effective progress-capturing predicates for diverse target states involves nontrivial reasoning across the entire codebase, spanning multiple functions and their associated data and control flows. Simple prompting techniques, such as in-context learning, chain-of-thought (CoT) prompting, retrieval-augmented generation, etc., can hardly support sophisticated analysis at the repository level. To this end, Locus features an agentic synthesizer-validator workflow, equipped with diverse program analysis tools to support the traversal of the control flow graph, tracking data dependencies, retrieving function calls, and symbolic execution. They serve as the agent’s action space, allowing the LLM to reason via CoT and act by calling them (Yao et al., 2023) to iteratively propose, refine, and validate candidate predicates.
@@ 2.1. Directed Fuzzing
Canary  Directed fuzzing aims to generate inputs that drive the program execution to reach predefined program states. In this paper, we represent these states with canaries (Hazimeh et al., 2020), which are considered reached when the corresponding canary condition is satisfied. Early works like AFLGo(Böhme et al., 2017) and Beacon(Huang et al., 2022) treat canaries primarily as the reachability to particular program points, i.e., specific line numbers in a code file. Recent works (Hazimeh et al., 2020; Weissberg et al., 2024) adopt assertion-based canaries, where predicates are explicitly inserted to check vulnerability-triggering conditions at runtime. Two examples of canaries are highlighted in yellow in Figure 1.
@@ 2.2. Motivating Example
A simplified CFG of this vulnerability is shown in Figure 1(a), where all the nodes highlighted in gray are equidistant to the canary node (highlighted in yellow), but only one node PLTE is in the path that can lead to the vulnerability (annotated as the green arrows). Based on this, fuzzers can only perform conservative pruning (annotated as dashed arrows) while omitting a large part of the paths that are irrelevant to the vulnerability, e.g., all solid black arrows. Capturing the progress towards reaching this canary requires specialized knowledge about libpng parsing logic. However, developing such specialized progress-capturing state representations requires manual efforts from experts and cannot easily generalize to different vulnerabilities and programs.
@@
Figure 1(b) shows an example of the input generator synthesized by LLM-based approaches that enforced some specific input constraints. However, this approach is inherently limited as program inputs can have sophisticated structures that cannot be easily enforced at the input level, e.g., the PNG file contains a compressed data chunk IDAT. Therefore, the constraints on the input generated by the LLM often reduce to generic input grammars, e.g., png_sig_cmp only trivially ensures the input is a valid PNG file. We also note that the synthesized generators cannot be easily checked to determine whether they can effectively help reach the canary. For example, png_write_frame_head constrains the generated input to a special PNG type APNG that is essentially impossible to reach the canary, but automatically checking this fact is challenging. Moreover, to constrain the input to satisfy certain properties necessary to reach the target, the generator sometimes needs to repeat the input processing logic presented in the execution path. Such repeated execution introduces additional overhead.
@@
Existing fuzzers struggle with such a vulnerability due to the complex parsing logic. Since PNG chunks may appear in arbitrary order and many are optional, the primary parsing routine implements a loop in png_read_info to iteratively process each chunk. Within this loop, multiple branches exist, and each is responsible for handling a specific chunk type, e.g., IHDR, IDAT, PLTE. This loop makes it possible to identify whether a given PNG file contains a PLTE chunk. However, since these branches are syntactically parallel and executed without a fixed order, they all appear equidistant from the vulnerability site (png_set_PLTE) in the control flow graph. As a result, the naive path distances cannot distinguish between them to prioritize one path over the other. This explains why, in Table 2, all existing fuzzers take a substantial amount of time to reach this target state in libpng.
Locus starts with generating a predicate to check whether the input PNG file contains a PLTE chunk at the caller of the canary function ( 1). This implies that if the PNG file does not contain a PLTE chunk, the execution can terminate, as it is impossible to reach the target state. While this predicate is semantically correct, i.e., the symbolic execution confirms there is no feasible path to satisfy the target state while violating the generated predicate (Section 3.4), and can help filter out non-palette-based PNG files, it is almost redundant as there is an existing one immediately after the generated check (if (color_type & MASK_PALETTE)), and thus the generated predicate could barely help the fuzzer.
To address this issue, Locus includes another refinement iteration (Algorithm 1) to propagate this predicate to a location closer to the program entry such that the infeasible execution can be terminated earlier. By traversing the call graph and retrieving the functions along the call chain, the refinement proposes a new location in function png_read_info right after the parsing loop ( 2). Such a propagated predicates will also be validated again to ensure it remains a strict precondition to reach the target states. At the next iteration of refinement, Locus proposes to propagate this predicate further inside the branch that parses the IDAT chunk ( 3) and creates a new helper variable found_plte. That is because the specification of the PNG file requires that the optional PLTE chunk must appear before the IDAT chunk. Therefore, by the time the parsing procedure reaches the IDAT chunk, we can already determine whether the input PNG file contains a PLTE chunk.
@@ 3.1. Formalization
We next illustrate how our agentic synthesizer-validator workflow can produce an admissible instrumented program $P^{\prime}$ that provides rich semantic feedback to the directed fuzzer to reach vulnerable states $S_{v}$.
@@ 3.2. Agent Toolset
Existing agentic practices for software analyses, such as bug fixing or fault localization, are often equipped with lightweight command-line tools to perform local reasoning(Paul Gauthier, 2024; Zhang et al., 2024c; Yang et al., 2024b). This is effective because the root cause of a bug is often spatially close to the observable failure, allowing the agent to retrieve relevant context from a narrow portion of the entire codebase. However, since our task formulation permits predicates to be synthesized at arbitrary program points, purely lexical tools cannot capture the necessary semantic relationships, and the agent therefore needs additional tools to understand program behaviors.
To equip the agent with such code reasoning capability at arbitrary program points, we provide a suite of tools that Locus can invoke to navigate the entire codebase and retrieve relevant code snippets. Particularly, in addition to common tools like code search and file listing, Locus integrates specialized tools for traversing program graphs, including call graphs and reference graphs. Through the call graph APIs, the synthesizer agent can retrieve function call relationships and reason the interprocedural control flow, while the reference graph API allows it to identify variable usages, pointer dereferences, and data access patterns across the codebase. Table 1 shows the complete toolset and the ratio of their actual usage in our experiments. Among them, graph traversal accounts for over a quarter of all API invocations.
It is important to note that these graphs are only provided as supplementary references to Locus, as they are all derived statically and may miss program behaviors such as dynamic dispatches and indirect calls. We observe that the LLM used in Locus is capable of leveraging its learned knowledge to bridge this gap. For example, in the case of vulnerability TIF002 (see Table 2), Locus successfully resolved the indirect function pointer tif->tif_decoderow to its concrete implementation PixarLogDecode.
@@
2:a target-conditional equivalent program $P^{\prime}$
@@ 3.3. Synthesis
Algorithm 1 elaborates on the workflow of Locus. It iteratively localizes candidate program points and generates the predicates (lines 4–13), then refines and validates them for both syntactic and semantic correctness (line 8). Once validated, the predicates are used to instrument the original program so it remains fuzzing-admissible (line 16).
@@
For each approximated characteristic, the synthesizer needs to identify an appropriate program point $l$ where the predicate can be expressed in terms of the variables and expressions in scope. However, constraining LLM to directly identify the exact program point is often too challenging and unreliable, so the initial stage only asks the synthesizer to select a candidate function rather than a precise program point. Given the reasoning context generated by the LLM so far, the synthesizer attempts to generate an initial predicate $\phi_{l}$ in the candidate function. Once $\phi_{l}$ passes the validation stage (Section 3.4), we refine it in the following iterations. The goal of such a refinement is to move the predicate to a closer program point to the program entry, while preserving the same semantic meaning as the original predicate, e.g., checking for the same characteristics. This allows the program to reject invalid inputs sooner in the execution, enabling the fuzzer to explore more valid mutations within the same time budget.